User Tools

Site Tools


lbaops:lbamar2007:v211aholog

Disk set Xraid SAM1-7

The log from experiment v211a has been copied to: ftp.atnf.csiro.au:/pub/people/vlbi/incoming/

ANTAB and UVFLG files have been copied to the same place.

Additional experiment notes:

Observers: Claire, Aidan, Brett, Warren later

Agilent: 4.2GHz

Centre frequency: 1650MHz

DAS profile: VSOP_HO.PRO

SMY 2nd LO = 502MHz

Channel 1: IFP#1-LO, 1634-1650MHz, RCP, Coherence tone 163.9MHz

Channel 2: IFP#1-HI, 1650-1666MHz, RCP, Coherence tone 165.5MHz

Channel 3: IFP#2-LO, 1634-1650MHz, LCP, Coherence tone 163.9MHz

Channel 4: IFP#2-HI, 1650-1666MHz, LCP, Coherence tone 165.5MHz

2007.083.20:50:24.03;“Weather: freezing!, clear skies, light northerly breeze

2007.083.22:13:26.90;”fine, calm and cool. 1% cloud cover

2007.084.00:45:25.23;“20% cloud cover, calm, fine

2007.084.03:00:20.28;”70% cloud cover, light breeze

2007.084.05:11:02.48;“90% cloud cover, light breeze

2007.084.07:02:17.59;”100% overcast, calm

2007.084.08:15:55.43;“95% overcast, calm

regards,

Brett

lbaops/lbamar2007/v211aholog.txt · Last modified: 2015/12/18 16:38 (external edit)